OpenStudy (anonymous):

Find the difference quotient,f(x+h)-f(x)/h, for the function and simplify it. g(x)=9x2+9x-14 A)18xh+9h+9h^2 B)18x+9 C)18x+9+9h D)9x+6+18h what am i suppose to do?

OpenStudy (perl):

simplify the expression [g( x + h) - g(x)] / h

OpenStudy (anonymous):

@perl can u come and help?

OpenStudy (anonymous):

@perl i cant wrap my head around the problem, can you explain little more?

OpenStudy (campbell_st):

well you are being asked to use 1st principals... so \[g(x) = 9x^2 + 9x - 14\] and \[g(x+h) = 9(x + h)^2 + 9(x + h) - 14\] does that make sense?

OpenStudy (anonymous):

i think i do

OpenStudy (anonymous):

am i suppose to distribute the 9 and get (9x+9h)^2 and 9x+9h - 14 and get as a whole 18xh^2+9x+9h-14

OpenStudy (campbell_st):

no 9(x + h)2 = 9(x^2 + 2xh + h^2) = 9x^2 + 18xh + 9h^2 9(x + h) 9x + 9h

OpenStudy (campbell_st):

oops and 9(x + h) = 9x + 9h collect you like terms and simplify

OpenStudy (anonymous):

9x^2+18xh+9h^2+9x+9h?

OpenStudy (campbell_st):

well g(x+h) - g(x) = 9x^2 + 18xh + 9h^2 + 9x + 9h - 14 - 9x^2 +9x + 14 ------------ ------------------------------------------------- h h simplify this